# ATLAS: Adaptive Self-Evolutionary Research Agent with Task-Distributed Multi-LLM Supporters

### v2

@@ Abstract
Recent multi-LLM agent systems have shown promising capabilities for prompt optimization and automated problem solving, but prior approaches rely on frozen fine-tuned agents or a single fine-tuning pipeline. We propose an ATLAS (Adaptive Task-distributed Learning for Agentic Self-evolution) model in which multiple supporter agents collaboratively train and refine a self-evolving agent toward a domain-specific research agent. Supporter agents specialize in complementary roles, including exploration support, fine-tuning strategy, and adaptive reference-policy inspection for the proposed Evolving Direct Preference Optimization (EvoDPO). Within this framework, we introduce a diverse set of challenging problem instances and demonstrate that our method achieves substantial performance improvements over existing approaches across long horizons.
Machine Learning, ICML
@@ 1 Introduction
Fine-tuned large language model (LLM) agents are broadly used as components in agentic systems for complex problem solving, scientific computing, and code-generation workflows. Recent frameworks show that collections of specialized agents, combined with experience banks, routing, or collaborative coordination, improve task performance and reliability by optimizing search, planning, decision-making, strategies, or data generation (Liu et al., 2025; Jiang and Karniadakis, 2025; Wu et al., 2025; He et al., 2025; Chen et al., 2025). In parallel, multi-agent infrastructure and benchmarking efforts have made it easier to orchestrate such systems and study their behaviors at scale (Chen et al., 2023; Wu et al., 2023; Chen et al., 2024; Huang et al., 2025; Becker et al., 2025).
@@
In this work, we propose ATLAS (Adaptive Task-distributed Learning for Agentic Self-evolution), a framework targeting specialized research agent formation via systematic model development. Rather than relying on a single static agent, we fine-tune a set of supporter LLM agents that collectively cultivate and refine a domain-specific research agent. The supporter agents are assigned complementary roles that support the learning dynamics: (i) exploration support to suggest diverse exploration strategies (ii) fine-tuning supervision to stabilize preference-based updates and prevent excessive deviation, and (iii) reference-policy inspection to decide when and how to renew the reference policy for well-aligned adaptation. By explicitly separating these tasks, the training loop prioritizes progressive agent improvement rather than merely search or planning enhancements, and it remains compatible with scalable multi-agent execution (Xie et al., 2024).
@@
Our learning mechanism leverages preference optimization as a practical alternative to Reinforcement Learning from Human Feedback (RLHF) pipelines. Direct Preference Optimization (DPO) enables preference learning without an explicit reward model (Rafailov et al., 2024), and recent work extends preference learning beyond simple binary comparison and introduces progressive schemes to handle distribution shift (Liao et al., 2025; Yang et al., 2025). Our work aligns with evolutionary or multi-agent fine-tuning approaches that iteratively improve a developing agent using critique or debate signals (Surina et al., 2025; Zhou et al., 2025). However, a key limitation in many iterative preference-based pipelines is fixed reference policies, which leads to misaligned references, overly conservative updates, or stagnation.
@@
To address this, we introduce Evolving DPO (EvoDPO), a preference-optimization loop with telemetry-driven fine-tuning control and adaptive reference management. At each fine-tuning phase, a strategist agent tunes DPO hyperparameters based on training diagnostics. In parallel, EvoDPO updates its policy using the DPO algorithm, while modeling the reference policy as a phase-indexed variable. Specifically, the self-evolving agent proposes a Kullback-Leibler (KL) regularized reference policy to be used in the next phase. Then, a policy-inspector agent applies a KL trust-region gate to decide whether to promote the proposed policy as the next reference or to keep the existing one. This integration of adaptive preference optimization and task-distributed agents prevents stagnation from stale references while avoiding instability from overly aggressive reference adaptation, enabling robust long-horizon self-evolution. The resulting mechanism is complementary to recent multi-agent preference-alignment approaches (Lyu et al., 2025).
To rigorously evaluate our core algorithm, EvoDPO, we conduct a comparative analysis against a fixed-reference DPO baseline within a theoretical framework. Specifically, we formulate a non-stationary contextual bandit problem and analyze the dynamic regret to formally characterize and demonstrate the performance advantages of our algorithm.
In the numerical results section, we compare our proposed methods against EvoTune, a single-agent evolution baseline that performs isolated evolutionary exploration and fixed-reference fine-tuning (Surina et al., 2025). We evaluate our framework on two complementary settings. First, we consider contextual multi-armed bandits under non-stationary environments (Wu et al., 2018), providing a controlled setting for adaptive policy evolution under the drifting model parameter and connecting to recent LLM-assisted bandit initialization ideas (Alamdari et al., 2024; Toscano et al., 2025). Second, to demonstrate applicability to scientific machine learning (SciML), we address the 1D Burgers’ equation using physics-informed neural networks (PINNs). Instead of relying on fixed, hand-tuned loss weights, ATLAS adaptively reweights PINN loss terms, effectively designing the training objective during optimization. This is particularly relevant in challenging regimes, such as those exhibiting near finite-time blow-up behavior (Kumar and Mukherjee, 2024). Overall, these evaluations indicate that our framework can sustain iterative improvement in both decision-making and SciML settings.
In this work, our contributions are as follows:
(i) Task-distributed multi-LLM agents structure for self-evolving research agent formation
(ii) Adaptive reference management mechanism for the EvoDPO algorithm
(iii) Numerical results showing optimal problem solving under long-horizon self-improvement
@@ 2 Proposed Methods
2.1 ATLAS : Adaptive Task-distributed Learning for Agentic Self-evolution
ATLAS is the end-to-end self-evolution framework comprising (i) a task-distributed multi-LLM supporter layer, (ii) a preference data construction procedure, and (iii) an adaptive fine-tuning loop for a developing research agent. Figure 1 illustrates the overall architecture of our framework and the specific workflow of each separated agent.
Self-evolving research agent.
Before each fine-tuning stage, the research agent performs a six-island exploration procedure in parallel to generate a diverse set of candidates (Surina et al., 2025). Each island maintains an independent candidate population and an update rule that mitigates mode collapse and encourages specialization across islands. For each explored candidate, the agent stores the task score and auxiliary metadata in an island-specific buffer. To preserve diversity, the agent also maintains a clustering over buffered candidates and tracks cluster coverage during selection. In particular, the agent monitors the number of clusters represented both in the buffer and within the count-based top-$s$ candidates, and the agent prioritizes candidates that preserve diverse coverage rather than concentrating on a single mode. At each exploration round, each island ranks candidates by score and uses the top-$s$ to form preference pairs. This procedure biases exploration toward high-scoring regions while preserving diversity via the clustering signal.
Exploration is implemented through a producer-consumer architecture. The producer assigns prompts to each island and the consumers evaluate candidates to produce task scores and metadata. The system also selects the inference model used by each island, enabling consistent generation behavior within an island while allowing diversity across islands.
For preference learning, the agent partitions evaluated candidates into passed and failed sets according to the task evaluation. It then selects count-based top-$s$ passed candidates to form preferred and less-preferred pairs $(y^{+},y^{-})$, and attaches auxiliary metadata $m$ for filtering and analysis. This process yields a specific preference dataset $D_{k}$
$D_{k}=\{(x_{i},y_{i}^{+},y_{i}^{-},m_{i})\}_{i=1}^{N_{k}},$
where $x$ denotes the prompt and $N_{k}$ is the number of preference pairs collected in phase $k$.
@@ 2.1 ATLAS : Adaptive Task-distributed Learning for Agentic Self-evolution
Figure 1: ATLAS workflow. ATLAS (Adaptive Task-distributed Learning for Agentic Self-evolution) alternates between (i) exploration with a supporter agent to generate diverse candidates and a preference dataset, and (ii) EvoDPO updates consisting of DPO fine-tuning (strategist-guided) and reference promotion via an inspector gate based on score improvement and a KL budget.
Let $\pi_{\theta,k}$ denote the current research agent policy and $\pi_{ref,k}$ the reference policy used at fine-tuning phase $k$. After constructing $D_{k}$, the research agent updates the policy by minimizing the DPO objective, and proposes a KL-regularized reference for the next phase $k+1$.
Task-Distributed Multi-LLM Supporters.
In our multi-LLM agentic system, the framework consists of (i) a developing Research Agent that is iteratively updated, and (ii) a set of frozen, fine-tuned supporter LLM agents specialized for distinct roles. We initialize the developing research agent from a general pre-trained model, whereas the supporter agents start from instruction-tuned or domain-adapted checkpoints. This asymmetric initialization enables the supporter agents to provide strong guidance and critique, particularly in early training.

### v3

# ATLAS: A Multi-LLM Training Framework for EvoDPO with Adaptive Reference Evolution

@@ Abstract
Recent multi-LLM agent systems have shown promising capabilities for automated problem-solving, yet they predominantly rely on frozen agents or static fine-tuning pipelines. To address this limitation, our primary contribution is ATLAS (Adaptive Task-distributed Learning for Agentic Self-evolution), a multi-agent framework where specialized meta-agents collaboratively train and refine an active agent toward a domain-specific policy. A core challenge in iterative preference learning within these pipelines is the reliance on fixed reference models, which typically leads to overly conservative updates or training stagnation. To overcome this, the framework’s algorithmic engine utilizes Evolving Direct Preference Optimization (EvoDPO). EvoDPO employs an inspection agent to perform adaptive, proxy-KL gated reference policy updates based on continuous training telemetry. We evaluate this full framework across a diverse set of challenging environments—including non-stationary contextual bandits, partial differential equations (PINNs), and combinatorial optimization tasks (TSP, Bin Packing). Through comparison against fixed-reference, adaptive-reference, and external automated-discovery baselines, our results suggest that ATLAS combines supporter-driven exploration with EvoDPO-driven stability to improve long-horizon evaluator-driven self-improvement.
@@ 1 Introduction
Fine-tuned large language model (LLM) agents are broadly used as components in agentic systems for complex problem solving, scientific computing, and code-generation workflows. Recent frameworks show that collections of specialized agents, combined with experience banks, routing, or collaborative coordination, improve task performance and reliability by optimizing search, planning, decision-making, strategies, or data generation Liu et al. (2025); Jiang and Karniadakis (2025); Wu et al. (2025); He et al. (2025); Chen et al. (2025). In parallel, multi-agent infrastructure and benchmarking efforts have made it easier to orchestrate such systems and study their behaviors at scale Chen et al. (2023); Wu et al. (2023); Chen et al. (2024); Huang et al. (2025); Becker et al. (2025).
@@
In this work, we propose ATLAS (Adaptive Task-distributed Learning for Agentic Self-evolution), a framework targeting specialized Evolving Agent formation via systematic model development. Rather than relying on a single static agent, we fine-tune a set of supporter LLM agents that collectively cultivate and refine a domain-specific Evolving Agent. The supporter agents are assigned complementary roles that support the learning dynamics: (i) exploration support to suggest diverse exploration strategies (ii) fine-tuning supervision to stabilize preference-based updates and prevent excessive deviation, and (iii) reference-policy inspection to decide when and how to renew the reference policy for well-aligned adaptation. By explicitly separating these tasks, the training loop prioritizes progressive agent improvement rather than merely search or planning enhancements, and it remains compatible with scalable multi-agent execution Xie et al. (2024).
@@
Our learning mechanism leverages preference optimization as a practical alternative to Reinforcement Learning from Human Feedback (RLHF) pipelines. Direct Preference Optimization (DPO) enables preference learning without an explicit reward model Rafailov et al. (2024), and recent work extends preference learning beyond simple binary comparison and introduces progressive schemes to handle distribution shift Liao et al. (2025); Yang et al. (2025). Our work aligns with evolutionary or multi-agent fine-tuning approaches that iteratively improve an Evolving Agent using critique or debate signals Surina et al. (2025); Zhou et al. (2025). However, a key limitation in many iterative preference-based pipelines is fixed reference policies, which leads to misaligned references, overly conservative updates, or stagnation.
@@
To address this, we introduce Evolving DPO (EvoDPO), a preference-optimization loop with telemetry-driven fine-tuning control and adaptive reference management. At each fine-tuning phase, a strategist agent tunes DPO hyperparameters based on training diagnostics. In parallel, EvoDPO updates its policy using the DPO algorithm, while modeling the reference policy as a phase-indexed variable. Specifically, the Evolving Agent proposes a Kullback-Leibler (KL) regularized reference policy to be used in the next phase. Then, a policy-inspector agent applies a proxy-KL gate to decide whether to promote the proposed policy as the next reference or to keep the existing one. This integration of adaptive preference optimization and task-distributed agents prevents stagnation from stale references while avoiding instability from overly aggressive reference adaptation, enabling robust long-horizon self-evolution. The resulting mechanism is complementary to recent multi-agent preference-alignment approaches Lyu et al. (2025).
We analyze EvoDPO through an idealized non-stationary preference contextual bandit, where the regret decomposition separates reference-induced bias from learning error under drift.
Empirically, we evaluate ATLAS on four executable optimization domains: non-stationary contextual bandits, PINN loss reweighting for Burgers’ equation, TSP, and Bin Packing. We compare against EvoTune Surina et al. (2025), AutoGen Wu et al. (2023), and FunSearch-style evaluator-driven program search.
Our contributions are as follows:
(i) We introduce ATLAS, a role-specialized multi-LLM training framework for evaluator-driven executable self-improvement. ATLAS separates candidate exploration, fine-tuning control, and reference-policy inspection rather than treating multi-agent interaction as a single monolithic search procedure.
(ii) We propose EvoDPO, an adaptive-reference variant of DPO that treats the reference policy as a phase-indexed variable and promotes a new reference only when it improves evaluator score while satisfying an empirical proxy-KL constraint.
(iii) We show that ATLAS gains arise from two complementary mechanisms: supporter agents improve the candidate distribution by expanding the search space, while EvoDPO stabilizes iterative preference optimization by mitigating stale-reference stagnation and uncontrolled reference drift.
(iv) We evaluate ATLAS across four executable optimization domains, including non-stationary bandits, PINN loss reweighting, TSP, and Bin Packing, and include ablations that isolate fixed-reference training, adaptive-reference training without supporters, and simplified reference-promotion alternatives.
@@ 2 Proposed Methods
2.1 ATLAS Architecture and Role-Specialization
We propose ATLAS as a collaborative, multi-agent hierarchy designed to decouple code exploration from optimization safety. Rather than relying on a single generalist model, we distribute distinct responsibilities across specialized LLMs. At the core of ATLAS is the Evolving Agent (Llama-3.2-1B), which is the only model updated by preference optimization. At phase $k$, the Evolving Agent receives prompts $x_{t}\sim\mathcal{X}$ and generates candidate executable programs or policies.
To construct $\mathcal{D}_{k}$, each candidate is evaluated by the corresponding task evaluator—simulating the non-stationary bandit, computing the PINN validation metric, or executing the heuristic solver for TSP and Bin Packing—to obtain a scalar utility score. We then form pairwise preferences by assigning the higher-scoring candidate as $y^{+}$ and the lower-scoring candidate as $y^{-}$. A pair is added to $\mathcal{D}_{k}$ only if its score gap exceeds a phase-dependent margin threshold. Importantly, supporter LLMs do not assign preference labels; they only propose candidate modifications or training-control suggestions, while all preference labels are induced by executable evaluator scores.
To guide this generation, an Exploration Supporter (gpt-oss-120B) performs static analysis on the Evolving Agent’s output, injecting domain-specific architectural patches (e.g., dynamic normalization, Huber loss) to prevent mode collapse. During the EvoDPO update, a Fine-Tuning Strategist (Qwen 3-32B) regulates the learning dynamics by monitoring phase-level training telemetry, such as score distributions, to dynamically modulate the DPO inverse-temperature $\beta_{\mathrm{DPO}}$ and pair-selection thresholds. Finally, to ensure stable progression, a Policy Inspector (Llama4-latest) operates as a strict safety gate for adaptive reference management. It evaluates the proposed KL-regularized policy $\pi^{kl}_{k}$ against the current reference $\pi_{\mathrm{ref},k}$ using an empirical trajectory-conditioned proxy $\widehat{\mathrm{KL}}_{k}$. This inspector enforces a proxy-KL gate, accepting updates only if the candidate demonstrates sufficient utility improvement ($\Delta\hat{S}_{k}\geq\epsilon_{s}$) without violating the predefined drift budget ($\delta_{H}$).
For full reproducibility, the exact system prompts, API configurations, and concrete examples of architectural patches generated by the supporter agents are detailed in Appendix C.
@@ 2.1 ATLAS Architecture and Role-Specialization
Figure 1: ATLAS workflow. ATLAS alternates between supporter-guided candidate exploration and EvoDPO updates with strategist-guided fine-tuning and proxy-KL-gated reference promotion.
2.2 Evolving DPO algorithm
Direct Preference Optimization.
Direct preference optimization (DPO) Rafailov et al. (2024) is an efficient alternative to reward-model-based RLHF. It optimizes a policy from pairwise preferences and avoids explicit reward learning and value-function estimation. Given a preference dataset $D$ of $(x,y^{+},y^{-})$, DPO minimizes
$L_{\mathrm{DPO}}$ $=-\mathbb{E}_{(x,y^{+},y^{-})\sim D}\Biggl[\log\sigma\Biggl(\beta_{DPO}\log\frac{\pi_{\theta}(y^{+}\mid x)}{\pi_{\mathrm{ref}}(y^{+}\mid x)}-\beta_{DPO}\log\frac{\pi_{\theta}(y^{-}\mid x)}{\pi_{\mathrm{ref}}(y^{-}\mid x)}\Biggr)\Biggr]$ (1)
where $\pi_{\mathrm{ref}}$ is a fixed reference policy and $\beta_{DPO}>0$ controls the strength of the preference margin.
A practical assumption behind (1) is that the preference data distribution is reasonably aligned with the reference policy: many public preference datasets are collected from, or near, a supervised fine-tuning (SFT) policy, and optimization is most stable when the likelihood ratios in (1) are not evaluated far off-distribution. In long-horizon phase-indexed loops, however, a fixed $\pi_{\mathrm{ref}}$ can become outdated as the Evolving Agent improves its policy and dataset. In this case, the mismatch between $D_{k}$ and $\pi_{\mathrm{ref}}$ can yield slow adaptation and misalignment.
Evolving DPO.
To address reference-data mismatch in long-horizon self-improvement, we introduce Evolving DPO (EvoDPO), which replaces the fixed reference with a phase-indexed reference $\pi_{\mathrm{ref},k}$ that updates across phases. At phase $k$, the Evolving Agent is updated by optimizing a DPO objective anchored at the current reference:
$L_{\mathrm{EvoDPO}}$ $=-\mathbb{E}_{(x,y^{+},y^{-})\sim D_{k}}\Biggl[\log\sigma\Biggl(\beta_{DPO}\log\frac{\pi_{\theta}(y^{+}\mid x)}{\pi_{\mathrm{ref},k}(y^{+}\mid x)}-\beta_{DPO}\log\frac{\pi_{\theta}(y^{-}\mid x)}{\pi_{\mathrm{ref},k}(y^{-}\mid x)}\Biggr)\Biggr]$ (2)
yielding an updated Evolving Agent policy $\pi_{\theta,k+1}$. The reference is then updated conservatively by the reference management.
Empirical KL estimate in the gate.
The exact prompt-conditional divergence $D_{KL}$
$D_{KL}\big(\pi(\cdot\mid x)\,\|\,\pi_{\mathrm{ref},k}(\cdot\mid x)\big)$
and the empirical gate statistic as
Inspector score functional.
At phase $k$, the Policy Inspector decides whether to promote the reference using a randomly sampled gate subset $G_{k}$ drawn from the current phase buffer. Each element of $G_{k}$ is a prompt $x$ together with its associated evaluated candidates and task scores collected during exploration. This design aligns the gate with the on-distribution prompts encountered by the Evolving Agent during self-evolution, while reducing selection adaptivity relative to score-based selection.
Reference-improvement operator.
In practice, we approximate the ideal reference-improvement operator by optimizing over a finite candidate set $\mathcal{C}_{k}\subset\Pi$. At phase $k$, we construct $\mathcal{C}_{k}$ from (i) a small number of intermediate checkpoints of the Evolving Agent collected during the phase, and (ii) candidate policies produced under the exploration and guidance from the Fine-Tuning Strategist. Thus, $\mathcal{C}_{k}=\{\pi^{(1)}_{k},\ldots,\pi^{(M)}_{k}\}$, and $\mathcal{C}_{k}$ always includes the phase-final policy $\pi_{\theta,k+1}$. With this candidate set, we define the practical reference-improvement operator as
$\pi^{kl}_{k}:=\arg\max_{\pi\in\mathcal{C}_{k}}\Bigl\{\widehat{\mathcal{S}}_{k}(\pi)-\beta_{\mathrm{ref}}\,\widehat{\mathrm{KL}}_{k}\bigl(\pi\,\|\,\pi_{\mathrm{ref},k}\bigr)\Bigr\},$ (4)
where $\widehat{\mathcal{S}}_{k}$ and $\widehat{\mathrm{KL}}_{k}$ are computed using the phase-$k$ gate subset. Here, $\beta_{\mathrm{ref}}>0$ is a reference-gating trade-off coefficient that controls how strongly the inspector penalizes reference drift during selection.
Algorithm 1 EvoDPO reference update
1:Initialize: Evolving Agent policy $\pi_{\theta,1}$, Reference $\pi_{\text{ref},1}\leftarrow\pi_{\theta,1}$
2:Hyperparameters: DPO temperature $\beta_{DPO}$, KL budget $\delta_{H}$, score tolerance $\epsilon_{s}$
3:for $k=1,\dots,K$ do
4:  Collect $D_{k}=\{(x_{i},y_{i}^{+},y_{i}^{-},m_{i})\}_{i=1}^{N_{k}}$
5:  Update Evolving Agent policy $\pi_{\theta,k+1}\leftarrow\arg\min_{\pi\in\Pi}\mathcal{L}_{\text{EvoDPO}}(\pi;\pi_{\text{ref},k},\mathcal{D}_{k})$
6:  Compute $\pi^{\{k\}}_{k}$ via Eq. (4)
7:  Compute $\widehat{\mathrm{KL}}_{k}(\pi^{\{k\}}_{k}\,\|\,\pi_{\mathrm{ref},k})$
8:  if $\Delta\widehat{\mathcal{S}}_{k}\geq\epsilon_{s}$ and $\widehat{\mathrm{KL}}_{k}(\pi^{\{k\}}_{k}\,\|\,\pi_{\mathrm{ref},k})\leq\delta_{H}$ then
9:   Accept $\pi_{\text{ref},k+1}\leftarrow\pi^{\{k\}}_{k}$
10:  else
11:   Reject $\pi_{\text{ref},k+1}\leftarrow\pi_{\text{ref},k}$
12:  end if
13:end for
Reference update with conservative acceptance.
At phase $k$, the Evolving Agent proposes the candidate reference $\pi^{kl}_{k}$ via Eq.  (4). Then the Policy Inspector applies a gating rule using (i) a score improvement $\Delta\widehat{\mathcal{S}}_{k}$ and (ii) a proxy-KL gate measured by the trajectory-conditioned token-level KL estimator $\widehat{\mathrm{KL}}_{k}$ on the phase-$k$ gate subset $G_{k}$. The system accepts a reference promotion if
2.3 Theorem for Adaptive Reference Analysis
Scope of the theory.
The regret analysis is intended to isolate the effect of adaptive reference progression under drift, not to provide an end-to-end guarantee for full sequence-level LLM fine-tuning. The theoretical model abstracts EvoDPO as a non-stationary preference contextual bandit with a KL-regularized comparator. Therefore, the theorem does not directly model the finite candidate set, the empirical token-level KL estimator, or the discrete accept/reject inspector gate used in implementation. These components are treated as practical approximations of the idealized reference-progression mechanism and are evaluated empirically through ablations, acceptance-rate statistics, and sensitivity analyses. The non-stationary contextual bandit experiment in this section is therefore best interpreted as a reward-feedback stress test of adaptive reference management, rather than a direct instantiation of the Bradley–Terry preference-bandit abstraction used in the theorem.
Problem Formulation
$u_{t}(x,y):=\langle\theta_{t},\phi(x,y)\rangle$ (6)
$\sum^{T}_{t=2}\|\theta_{t}-\theta_{t-1}\|_{2}\leq V_{T}$ (8)
Define the value of the policy at time $t$ as
$\pi_{t}:=\arg\max_{\pi\in\Pi}\left\{\widehat{J}_{t}(\pi)-\beta_{\mathrm{ref}}D_{\mathrm{KL}}\bigl(\pi(\cdot|x_{t})\|\pi_{\mathrm{ref},t-1}(\cdot|x_{t})\bigr)\right\}$ (10)
Here $\widehat{J}_{t}$ is induced by the sliding-window preference estimator in Appendix B; this idealized policy is used only for the regret analysis and should not be interpreted as the exact finite-sample sequence-level DPO update used in implementation.
$\pi^{*}_{t}:=\arg\max_{\pi\in\Pi}J_{t}(\pi)$ (11)
$\pi^{kl}_{t}:=\arg\max_{\pi\in\Pi}\bigl\{J_{t}(\pi)-\beta_{\mathrm{ref}}D_{KL}(\pi(\cdot|x_{t})||\pi_{ref,t-1}(\cdot|x_{t}))\bigr\}$ (12)
We define cumulative dynamic regret as
$R_{T}=\sum^{T}_{t=1}r_{t}.$ (13)
Dynamic Regret Analysis.
The cumulative dynamic regret decomposes as
$R_{T}=R_{T}^{\mathrm{error}}+R_{T}^{\mathrm{bias}}.$
Combining the learning-error bound with the reference-induced bias bound gives, with probability at least $1-\delta$,
$R_{T}\leq\mathcal{O}\left(T^{\kappa}V_{T}+T^{1-\frac{\kappa}{2}}\sqrt{\log(T/\delta)}+T^{1-\kappa}+V_{T}\right),$ (14)
where the initial reference-mismatch term is absorbed into the constant. The formal derivation is provided in Appendix B.
3 Numerical Results
In this section, we conduct a comprehensive set of experiments to evaluate our proposed methods. The main finding is that structured collaboration between an Evolving Agent and Role-Specialized Multi-LLM Training supporters enables the ATLAS to solve complex optimization problems that require adaptation to non-stationary environments and time-dependent PDE across varying Reynolds number.
3.1 Experimental Setup and Domains
We evaluate ATLAS’s capacity for long-horizon self-improvement across diverse optimization landscapes. We deploy the framework in four challenging domains. First, to test decision-making under concept drift, we evaluate the system on a non-stationary contextual bandit, where the underlying environment parameter shifts continuously under a known total variation budget. Second, to demonstrate continuous optimization capabilities in scientific machine learning (SciML), we task the agent with solving the 1D viscous Burgers’ equation using a PINN. Here, the Evolving Agent must adaptively reweight a composite loss function to prevent mode collapse as a drifting viscosity schedule pushes the environment from smooth diffusion to a stiff, convection-dominated shock regime. Finally, to assess discrete combinatorial generalization, we introduce Traveling Salesperson Problem (TSP) and Bin Packing tasks, requiring the agent to design scalable heuristic routing policies. Across all domains, we benchmark ATLAS against fixed-reference evolution (EvoTune) and representative automated-discovery baselines, including FunSearch and AutoGen. Full mathematical formulations and hyperparameter configurations for all environments are detailed in Appendix C.1.
Baselines.
We evaluate the performance of our full framework, ATLAS, through an ablation study and a comparison against a standard baseline (detailed in Appendix C.1) to isolate the contributions of our proposed components:
(i) EvoTune: A baseline method that utilizes a fixed reference policy and maintains static hyperparameters over the entire course of training.
(ii) FunSearch: We implement FunSearch as an evaluator-driven code-generation baseline that iteratively proposes candidate programs, scores them using the same task evaluator, and retains high-performing programs in an experience buffer. Unlike ATLAS, FunSearch does not update the underlying Evolving Agent with DPO and does not use an adaptive reference model.
(iii) AutoGen: We implement AutoGen as a multi-agent conversation baseline in which agents propose, critique, and revise executable candidate solutions using the same task-specific evaluator. Unlike ATLAS, AutoGen performs inference-time coordination only and does not fine-tune the Evolving Agent or apply proxy-KL gate reference promotion.
3.2 Results
(a) Bandit
(b) PINN
(c) TSP
(d) Bin Packing
Figure 2: Experimental results across four executable optimization domains. (a) Bandit Negative Mean Regret (NMR), shown with a truncated y-axis for readability. (b) PINN fixed-evaluator validation loss on a log scale. (c) TSP evaluator score. (d) Bin Packing evaluator score. Shaded regions denote the standard error of the mean (SEM) across 10 independent seeds.
Table 1: Main 10-seed performance summary across four executable optimization domains. For Bandit, TSP, and Bin Packing, values report relative improvement from the initial policy to the final policy, where higher is better. For PINN, values report the final fixed-evaluator validation loss, where lower is better.
Method Bandit Improve. $\uparrow$ PINN Final Loss $\downarrow$ TSP Improve. $\uparrow$ Bin Packing Improve. $\uparrow$
EvoTune $20.23\%$ $1.80\times 10^{-3}$ $58\% 2.442.$$4\%$
AutoGen $2.78\%$ $3.214\times 10^{-1}$ $10.17\%$ $4.36\%$
FunSearch $20.16\%$ $4.70\times 10^{-4}$ $20.36\%$ $2.10\%$
ATLAS $\mathbf{20.71\%}$ $\mathbf{7.70\times 10^{-5}}$ $\mathbf{25.52\%}$ $\mathbf{5.06\%}$
Experimental setup.
We evaluate ATLAS across four executable optimization domains: non-stationary contextual bandits, PINN loss reweighting for Burgers’ equation, Traveling Salesperson Problem (TSP), and Bin Packing. Figure 2 shows the learning trajectories over self-improvement rounds, and Table 1 summarizes final performance.
Each task is evaluated using its own task-specific metric. For Bandit, we report relative improvement in Negative Mean Regret (NMR). For PINN, we report the final fixed-evaluator validation loss. For TSP, we report relative improvement in evaluator score based on the negative optimality gap against the exact-solver reference. For Bin Packing, we report relative improvement in evaluator score based on the negative extra-bin count relative to the integer-programming reference. Full hyperparameters, hardware details, seed counts, and task configurations are provided in Appendix C.5.
Bandit.
In the non-stationary contextual bandit task, ATLAS achieves the largest relative improvement in Negative Mean Regret (NMR). ATLAS improves by $20.71\%$, compared with $20.23\%$ for EvoTune, $20.16\%$ for FunSearch, and $2.78\%$ for AutoGen. Although the margin over EvoTune and FunSearch is modest, ATLAS maintains the best final trajectory in Figure 2(a), suggesting that supporter-guided exploration and adaptive reference management provide stable gains under parameter drift.
PINN.
For the PINN task, candidate-generated losses are used only for training. All candidates are ranked and reported using a fixed external evaluator that is not generated by the agent, including normalized PDE residual, initial-condition error, boundary-condition error, and solution $L_{2}$ error against a high-resolution reference solution. This prevents trivial improvements caused by rescaling the generated training objective. ATLAS obtains the lowest final fixed-evaluator validation loss, $7.70\times 10^{-5}$, compared with $1.80\times 10^{-3}$ for EvoTune, $4.70\times 10^{-4}$ for FunSearch, and $3.214\times 10^{-1}$ for AutoGen. This indicates that ATLAS achieves stronger error reduction under the fixed evaluator while avoiding dependence on the generated training objective itself.
TSP.
For the Traveling Salesperson Problem, ATLAS improves the evaluator score by $25.52\%$, compared with $24.58\%$ for EvoTune, $20.36\%$ for FunSearch, and $10.17\%$ for AutoGen. The TSP evaluator measures tour quality using the negative optimality gap relative to an exact-solver reference, while also enforcing validity constraints such as visiting every node exactly once and returning to the start node. These results suggest that supporter-guided structural modifications, such as route repair and local-improvement heuristics, help the Evolving Agent move beyond simple greedy tour construction.
Bin Packing.
For Bin Packing, ATLAS improves the evaluator score by $5.06\%$, compared with $2.44\%$ for EvoTune, $2.10\%$ for FunSearch, and $4.36\%$ for AutoGen. The Bin Packing evaluator measures solution quality using the negative extra-bin count relative to an integer-programming reference, while checking feasibility constraints such as assigning every item exactly once and respecting bin capacities. The stronger improvement over EvoTune and FunSearch indicates that the supporter layer helps discover more effective packing heuristics, while the Policy Inspector prevents unstable reference promotions that improve isolated instances but harm feasibility or generalization.
3.3 Ablation Study
To isolate the contribution of each component, we evaluate the system incrementally in Table 2. Moving from EvoTune to EvoDPO isolates the effect of adaptive-reference preference optimization without the multi-LLM supporter layer. This transition substantially improves PINN performance, reducing validation loss from $1.829\times 10^{-3}$ to $3.53\times 10^{-4}$, improves TSP score from 262.8154-262.8154 to 261.7880-261.7880, and improves Bin Packing score from 2023.6720-2023.6720 to 2004.7350-2004.7350. These gains indicate that adaptive reference evolution can improve stability and late-phase adaptation in several domains even without external supporter agents.
Adding the role-specialized supporter layer yields the full ATLAS system and produces the best performance across all four tasks. ATLAS improves Bandit NMR to 0.4993-0.4993, reduces PINN validation loss to $7.7\times 10^{-5}$, improves TSP score to 259.4793-259.4793, and improves Bin Packing score to 1969.3510-1969.3510. These results support the interpretation that EvoDPO contributes reference-adaptive training stability, while supporter agents improve candidate quality and search diversity. The combination is especially effective in combinatorial tasks, where structural heuristic modifications help the Evolving Policy move beyond narrow candidate-generation patterns.
Table 2: Ablation study comparing final performance across all tasks. Bandit values report Negative Mean Regret (NMR), where higher is better. PINN values report fixed-evaluator validation loss, where lower is better. TSP and Bin Packing values report task-specific executable solver scores, where higher is better. Metrics are averaged across 10 independent seeds.
Configuration Bandit $\uparrow$ PINN $\downarrow$ TSP $\uparrow$ Bin Packing $\uparrow$
Baseline (EvoTune) -0.5023 0.001829 -262.8154 -2023.6720
+ Adaptivity (EvoDPO) -0.5023 0.000353 -261.7880 -2004.7350
+ LLM Supporters (ATLAS) -0.4993 0.000077 -259.4793 -1969.3510
The ablation results show that ATLAS gains come from complementary mechanisms. EvoDPO improves stability by reducing stale-reference effects under drifting preference data, while the supporter layer improves exploration by expanding the candidate distribution with structural edits, hyperparameter suggestions, and task-specific program modifications. Thus, ATLAS should be interpreted as combining supporter-driven exploration with gated adaptive-reference stabilization: EvoDPO does not replace exploration, but makes iterative fine-tuning less stagnant once stronger candidates are available.
4 Conclusion
We introduced ATLAS, a role-specialized multi-LLM training framework that decomposes long-horizon executable self-improvement into candidate exploration, fine-tuning control, and reference-policy inspection. The core algorithmic component, EvoDPO, performs preference optimization with adaptive reference evolution, promoting a new reference only when it improves evaluator score while satisfying an empirical proxy-KL gate constraint.
We analyzed an idealized non-stationary preference contextual bandit to isolate the benefit of adaptive reference progression under drift, rather than to claim an end-to-end guarantee for full sequence-level LLM fine-tuning. Empirically, we evaluated the implemented system across decision-making, scientific machine learning, and combinatorial optimization tasks. The results suggest that gated reference evolution can mitigate stale-reference stagnation while maintaining stable long-horizon improvement.
Future work will extend this framework to broader scientific discovery settings where evaluator signals are sparse, noisy, or expensive to obtain.
Appendix A ATLAS appendix
This appendix provides the full system-level execution procedure for ATLAS. While the main text focuses on the core EvoDPO update and the reference-promotion rule, Algorithm 2 describes how candidate generation, evaluator-based scoring, preference-pair construction, supporter feedback, and EvoDPO fine-tuning are combined into one long-horizon self-improvement loop.